### v3

# DrawTalking: Building Interactive Worlds by Sketching and Speaking

Abstract.
We introduce DrawTalking, an approach to building and controlling interactive worlds by sketching and speaking. It emphasizes user control and flexibility, and gives programming-like capability without requiring code. We built a prototype to demonstrate it.
An early open-ended study shows the mechanics resonate and are applicable to many creative-exploratory use cases, with the potential to inspire and inform research in future natural interfaces for creative exploration and authoring.
creativity, sketching, play, programmability, multimodal, human-AI collaboration, prototyping
ccs: Human-centered computingccs: Human-centered computing Graphical user interfacesccs: Human-centered computing Natural language interfacesccs: Human-centered computing Interaction techniquesccs: Human-centered computing Interactive systems and tools
Figure 1. Our approach DrawTalking mediates sketching and talking-out-loud through direct manipulation, enabling many use cases across improvisational creative tasks.
[…]
[Controls by sketching and speaking]This figure showcases the DrawTalking sketching + speaking controls. (1) On the left is a real-world photo of the user labeling a sketch on the digital canvas. The canvas contains a few user-drawn rectangle objects. The user narrates ”and I’m also going to have a rectangle,” which shows-up in a text transcript. Simultaneously, the user holds one of the rectangles with their finger and tapping the word ”rectangle” in the aforementioned transcript. This labels the user’s sketch with the word ”rectangle.” This is important because now the system knows what the name user’s object is, meaning it can now be controlled with more speech commands. (2) On the right is a mock-up of an example scenario; 3 panels. Panel 1: 3 lily pads have been drawn, as well as a frog. A caption reads a speech command: ”Forever this frog hops to a lily.” The user’s finger points to the frog to indicate which frog is being commanded. Dotted arrows indicate that the frog will now hop between the lilies once the command stops. Panel 2: Now a butterfly and some flowers have been drawn. The new command caption reads, ”The butterfly follows the frog.” Panel 3: The command caption reads ”This butterfly is slow” and the user’s finger is touching the butterfly, meaning the butterfly’s speed will now be decreased. Dotted arrows are now indicating that the butterfly has been moving around, following the frog as previously commanded. Slowing the butterfly does not require redoing the previous command.
1. Introduction
Sketching while speaking aids innovation, thinking, and communication — with applications in animation, game design, education, engineering, rapid prototyping, storytelling, and many other creative and spontaneous activities (Fan et al., 2023; Tversky, 2011). The combination enables us to think about and share anything through make-believe – including things that do not or cannot exist. We achieve this by assigning representations (sketches) to semantic concepts (objects, behaviors, properties) (Turner, 2016). For example we might suspend our disbelief so a square represents anything from a house, to a playing card, a map, dog, or person.
This work is an attempt to realize a style of spontaneous interaction that seamlessly integrates sketching and talking-out-loud to build interactive, explorable worlds and simulations, thus greatly increasing our range of computational expression (Victor, 2014). Prior works in interactive sketching (Sutherland, 1963; Kazi et al., 2014a; Suzuki et al., 2020; Perlin et al., 2018; Saquib et al., 2021), language/AI-mediated interfaces (Winograd, 1972; Bolt, 1980; Liu et al., 2023; Xia et al., 2023; Huang et al., 2020; Xia, 2020; Subramonyam et al., 2018; Laput et al., 2013), and visual programming-adjacent interfaces or games (Sutherland, 1966; Resnick et al., 2009; Landay, 1996; Perlin and Goldberg, 1996; cycling74, 2018; Little Big Planet 2, 2011; Dreams, 2020) has laid valuable groundwork. However, they often require content to be pre-built or assume that users have predetermined goals, limiting the user’s potential to be spontaneous. Tools focus on generating a specific output rather than facilitating an ongoing creative process (Compton and Mateas, 2015). They might enforce given representations of objects (e.g. realistic or specific sketch recognition). They might feature complex UI, and in the case of programming-oriented tools, require explicit programming knowledge (whether via text, nodes, or blocks).
In our prototype system, DrawTalking, users speak while freehand-sketching to create, control, and iterate on interactive visual mechanisms, simulations, and animations. Via speech and direct manipulation, the user names their sketches to provide semantic information to the system, and narrates desired behaviors and rules to influence the interactive simulation, as if by explaining to others or telling a story. This is inspired by the way in which people might speak and/or write text to describe objects as they are being drawn. The key insight is that in the context of narration, the machine can use this information directly without guessing-at the user’s intended meaning behind their drawings. As a result, users can build and control interactive worlds by simple combinations of sketching and speaking. DrawTalking requires no preparation or staging, and supports the user in making changes to content and behavior anytime. By design, we balance AI-automation with user-direction such that the user is the one who chooses how to represent content, when to do operations, and what logic to define.
In sum, we contribute:
  • DrawTalking: a novel sketching+speech interaction that balances user-agency with machine automation.
  • A demonstration of DrawTalking as a multi-touch prototype application for the iPad.
  • An early qualitative study of DrawTalking that reveals its use for emergent creativity and playful ideation
2. Related Work
Our research is based on a confluence of advances in multimodal, sketching, and programming interfaces.
2.1. Natural Language-Adjacent Interfaces
Systems such as SHRDLU (Winograd, 1972) and Put That There (Bolt, 1980) pioneered the vision of employing natural language to communicate with computers. Due to recent advances in speech recognition and natural language understanding, the popularity of this interaction modality has exploded, and has been used in a wide range of domains. For example, VoiceCut (Kim et al., 2019) and PixelTone (Laput et al., 2013) allow users to speak short phrases or sentences to perform desired operations in image editing applications, but these applications are heavily domain-specific. Tools like WordsEye (Coyne and Sproat, 2001), Scones (Huang et al., 2020), and CrossPower (Xia, 2020) enable scene generation or content editing via language, and interfaces such as Visual Captions (Liu et al., 2023), RealityTalk (Liao et al., 2022), and CrossTalk (Xia et al., 2023) use language structure to make content appear during talks or conversations. However, many of these interfaces tend to assume that the user knows what they want to create in advance — i.e. an end-product with an initial goal. They require an initial phase in which the user must define content and behavior up-front. This could limit open-ended exploration during the creative process, when the user does not necessary have an end-goal in mind. Further, the majority of such interfaces use language input to generate or spawn content without the user in the loop. An alternative is to enable greater interactive control and definition of the behavior of the content. Our approach emphasizes flexibility and user-control during the creative process. The user can define and iterate on content at any point. Our prototype specifically supports definition of behaviors within an interactive simulation. Within this prototype, we explore language input, combined with direct manipulation, as a way of empowering the user to program and control scenes interactively.
2.2. Dynamic Sketching Interfaces
HCI researchers have extensively explored sketching interfaces for dynamic and interactive visualizations ever since the first graphical user interface (GUI) SketchPad (Sutherland, 1963) and William Sutherland’s thesis, the forerunner of the visual programming language (Sutherland, 1966). Many works use direct manipulation and sketching techniques to help users craft interactive behaviors and toolsets for illustrated animation, UI, and visual-oriented programs. For example, works by Kazi et al. (2014b; 2014a), Landay et al. (1996), Saquib et al. (2021), and Jacobs et al. (Jacobs et al., 2018) focused on mixing illustration, programming, and prototyping. Programming by-demonstration is featured in works such as K-Sketch (Davis et al., 2008) and Rapido (Leiva et al., 2021). Scratch (Resnick et al., 2009) is a well-known visual programming environment mixing game-like interactions with user-provided content and images for a playful experience. (Subramonyam et al., 2020) supports the user in forming connections between texts and concepts to learn via active diagramming. Our interface is complementarily framed around correspondences between sketches and language elements, but we use sketch-language mapping as a control mechanism, enabling the user to create interactive simulations and behaviors for open-ended exploration.
Prior work also explored supporting development of
pre-programmed simulations and domain specific behaviors to craft interactive diagrams. In Chalktalk, for example, the system uses sketch recognition to map a user’s hand drawn sketches into corresponding dynamic, pre-programmed behaviors/visualizations (Perlin et al., 2018). More domain-specific tools like MathPad2 (LaViola and Zeleznik, 2004), Eddie (Sarracino et al., 2017), PhysInk (Scott and Davis, 2013) and SketchStory (Lee et al., 2013) use hand-drawn sketches and direct manipulation interactions to create interactive simulations in physics, math, and data visualization.
2.3. Programming-Like Interfaces
The customizability and flexibility of such a more general interface implies a need for programmability. Examples of these include real-time world simulation systems and programmable environments such as the SmallTalk programming language(Goldberg and Robson, 1983), Scratch (Resnick et al., 2009; Maloney et al., 2010) (and similar AI voice-enabled explorations like StoryCoder (Dietz et al., 2023)), Improv (Perlin and Goldberg, 1996), ChalkTalk(Perlin et al., 2018), and creative world-building games like the Little Big Planet series (Little Big Planet, 2008; Little Big Planet 2, 2011; Ross et al., 2012) and Dreams (Dreams, 2020). These encourage interactive building of scenes, games and stories. They combine with elements of interactive visual programming and drawing/sculpting with many types of content (2D, 2.5D, 3D, images). But all use explicit interfaces for programming or programming-like functionality (nodes, wires, text). We depart from explicit UI for sketching+programming-like capability, and largely replace many programming-like functionality with the use of language. Our direction explores the use of verbal, descriptive story narration together with other input modalities (i.e., touch and pen input) to create animated and interactive graphics through sketching.
3. Formative Steps
We derive a set of design goals by examining existing practices. This would also inform the development of interaction techniques and a prototype interface.
3.1. Methodology
We used a mixed-method approach for our study consisting of an analysis of online instructional videos and conducting a set of sketching design sessions.
To get a sense of the relationships between speech, text, and drawings, and to find examples of content that people created, we conducted an informal (non-exhaustive) search for example content online. We looked for examples that involved creating (or showing) rough sketches while speaking, e.g. from popular video channels and educational course recordings. (Refer to appendix Figure 14 for additional sample materials.)
To gain further insights about workflow and technical requirements, we conducted informal exercises with 6 participants P1init-P6init to observe casual sketching and speaking process with little to no preparation. All participants had some experience with sketching-out ideas (e.g. concepts, storyboarding, project/presentation sketching for game design). Each participant was asked to think of (at least) 1 personal topic they would feel comfortable narrating while drawing with freehand sketches. Participants used their choice of tools during the session (eg, MS Paint, Notability, basic tools in Photoshop), but we only allowed basic color selection and transformations to keep tool usage roughly the same. (Refer to appendix Figure 15 for P1’s result.)
3.2. Results / Observations
3.2.1. Association between spoken language, text, and drawing:
Participants used a variety of visual styles (abstract or cartoon), symbols, and diagrammatic elements to express their ideas. They would sometimes label their objects with names or use in-line text to describe elements in the screen for them to reference later (Agrawala et al., 2011). Pointing or proximity to sketches while using deictics also acts as a way to refer to objects (as observed in content recorded with cursors or speakers’ hands.) In short, semantics are associated with sketches in many ways, including through narration, or explicit text labels. The natural association between spoken language, text, and sketches inspired our our main interaction technique.
3.2.2. Temporal synchronization between speech and drawing:
In the content search and exercises, people do not tend to synchronize their speech exactly when drawing specific objects (or if they do, the delay between actions is unreliable). However the order in which people mention sketches verbally usually will correspond to pointing and sketching actions(Oviatt, 1999). This means we can use this ordering property with respect to speech to let the user map semantics to objects in sequence, but temporal synchronization is unreliable.
3.2.3. Drawn versus non-drawn spoken content:
Although narration was used to explain the content on-screen, not all content (such as additional animations in the case of produced-content or additional details) and not all narration mapped to each other. In other words, there is content that the user doesn’t necessarily describe via language and some content might be considered unnecessary to represent. Hence, both modalities are tightly intertwined to represent the complete picture.
3.2.4. Speed and flow:
People operate at many different speeds when drawing and talking, and unpredictably move back and forth between sketches as they’re iterated on. This means that we can’t force people to use a specific speed or cadence.
3.2.5. Content modification:
For the exercises, a participant would usually multitask and refine their sketches over time after initially mentioning the entity. For example, P1init added colors to their gray bird and pond sketches as they described their experience at the scene. Additionally, the participant might make verbal corrections ( e.g. P1init initially called a ”pond” a ”lake,” but self-corrected verbally). This means that speech errors are natural in real-life and we ought to provide ways to correct it (Suhm et al., 2001). As users’ intent change, they should be able to update their content.
3.2.6. Hierarchical object model:
When referring to objects on-screen, narrators would name objects, refer to existing objects to draw attention to them, and define hierarchies and relationships between objects. In short, sketched presentations and content implicitly encoded a hierarchical object model, describing entities and the relationships between them, as well as ways to refer back to them.
Overall, based on our formative observations, we needed to prioritize user control over a variety of possible creative workflows. Rather than a fully-automatic solution, we needed a synthesis of user-directed input and system-feedback in support of spontaneous creative processes.
3.3. Design Goals
In brief, we envision an interface for creative exploration with interactive capability that (a) does not impose many assumptions about the user’s intent or content, (b) focuses on the process, not just on the artifact, and (c) does not require programming knowledge. Above all, the user should have control. To that end, we wanted an interface that:
D1 has minimal system assumptions
where the user controls the creative process, and the representation and behavior for sketches
D2 is flexible, mutable, fluid, and playful
in that it supports improvisation, quick changes, and rapid iteration of ideas with a spontaneous feel, where operations are easily accessible and doable in any order
D3 is transparent and error tolerant
by telegraphing what the system’s understanding is and providing multiple opportunities for users to make changes or recover from system or user error.
D4 supports programming-like capability
without the need for a coding interface
4. DrawTalking
Figure 2. Interface Overview: Shown is an interface screen (from P4 in section 5). The toolbar (left) enables edit operations, e.g. copy, delete, attach/detach, save sketch. The transcript view displays the user’s speech input in an interactive panel, and the semantic diagram displays the machine’s understanding of the input. ”Speech controls” stage/confirm an action, discard input, and toggle speech recognition. ”Transcript controls” offer quick transcript text selection. The status bar displays the current color, a compass pointing ”up,” and has the pen/eraser state-change buttons and indicators. The scene shown is just before the user confirms a command for the utterance ”The character jumps on the platforms”; it selects the sketch labeled ”character” and all sketches labeled ”platform.” Tapping the ”language action” button (top-right) will stage the command and display the semantic diagram (top-left) representing the machine’s understanding of the input with selected objects displayed under their respective words; the user confirms the command by tapping again. This causes the character to jump on all of the platforms. For details on the workflow and commands, see subsection 4.2.
Figure 3. Overall workflow: Dog and boy’s infinite game of fetch.
a) From left to right, the user draws and labels them using multiple approaches at different stages of drawing. a.1) The user is midway through drawing the boy, but can label it using ”This is a boy” as the pen is interacting with the object. a.2) The ball is already drawn, but unlabeled; during the same sentence the user can quickly tap the ball to label it. a.3) The user draws some water and uses free-form speech to say ”water” without deixis (such as ”this”/”that”). The user can select the object with touch and simultaneously tap the word ”water” in the transcript label the object with the word in the transcript. a.4) Touch+pen on a word will remove the label. Adjectives work in the same way.
b) Left: Labeled sketches are commanded. Right: Interactive user-participation. The user can move objects as the system is simulating their movements, which will dynamically adjust as the user plays-around spontaneously.
[…]
[A mock-up of sketch characters and props playing fetch, and the user interacting by moving the characters around]A mock-up of sketch characters and props playing fetch, and the user interacting by moving the characters around. Panels 1-2 show the boy repeatedly throwing a ball into a pool of water, with the dog giving the ball to the boy afterwards, in an infinite cycle. Below the command reads, ”Over and over the boy throws the ball into the water and then the dog gives the ball to the boy.” Panels 3-4 show the user moving the boy sketch around as the command continues, forcing the dog to change its trajectory as the targets change during the simulation. It’s a demonstration of user-in-the-loop interaction.
We designed and developed DrawTalking, an approach to creating interactive worlds using freehand-sketching and speech. It is based on design goals derived from our formative study and implemented as a pen+multi-touch application.
4.1. Concept
In DrawTalking, speaking serves a dual-purpose: the user can explain concepts and tell stories, and at the same time refer to and label the objects in their sketched world with semantics (e.g. nouns, adjectives, adverbs) determining their names and behaviors.
The user-specified labeling tells the system what the objects are, irrespective of their visual representation, making them controllable via the narration.
The user can furthermore create the rules to automate behavior between objects in the simulation. Touch controls also allow the user to interact directly with the simulated world. This results in a free-form sandbox for animation and programming-like behavior that mixes direct user-control with machine automation. An overall workflow of DrawTalking in shown in Figure 3.
4.2. User Interface Components
The interface (Figure 2) exposes a transcript view (Figure 2) and semantics diagram (Figure 4) to make system understanding of input transparent, quickly editable and accessible, and error-robust, as per our interface design goals (section subsection 3.3).
Speech recognition is continuous for interactive use, so the transcript view lets the user visualize and optionally edit speech input (see Figure 5), assign names and attributes to sketches using the pen linking modality, and stage/confirm commands.
When ready, the user taps ”language action” to stage a command. The diagram appears and displays a visual for the machine’s understanding of the input. It provides a way to reassign objects within the command (if the user wants). The user confirms with the same button to execute the command, or cancels with the discard button.
Complementary to the sketching and language workflows, the find panel along with the transcript view enables the user to find and teleport to objects by semantic search, which helps avoid losing track of objects. Additionally, it enables a means to view, toggle, or delete the current commands and rules acting on the world simulation (See Figure 6).
In sum, all three of these sketching-language interface components combine to support the user’s process via our design goals. Together, they provide transparency into the system’s state, multiple forms of editability and error-tolerance, and multiple optional levels for user control.
4.3. Sketching & Labeling
Sketches are independently-movable freehand drawings, text, or numbers created by the user. The user labels their sketches to make them controllable: nouns (names) for unique identification and adjectives and adverbs (properties) for modulating sketch behavior. For flexibility, we offer two direct ways to label (or unlabel).
  1. (1)
    tap 1 to many sketches and speak with deixis (Stapleton, 2017) (e.g. ”this/that is a ¡noun¿”, ”this/those are ¡noun¿s”).
  2. (2)
    touch sketches + pen-tap words to link at any time with any text in the history, as opposed to only the current speech, enabling freer narration.
Labeling works in tandem with the user interface elements (4.2). See Figure 3 for an example workflow.
Figure 4. Semantics Diagram: An example workflow of error-correction (changing the target ”building” from red house to yellow skyscraper). (a) spoken user input in the transcript panel. (b) The semantics diagram generated from the user’s input to visualize system’s interpretation. (c-d) Pen+touch interaction between diagram nouns and objects in the scene (including unlabeled ones) allows for re-linking to modify or correct the machine’s selections. (e) Confirm language command. If a verb is unknown for a command, the user can pick from an auto-presented list of similar verbs, or otherwise cancel. In our implementation, if a sentence produces too long a diagram to fit in the window, the user can zoom-out independently from the canvas.
Figure 5. transcript view: Selection / Deselection of Words: Toggle off/on words via touch-dragging to modify input for the next command. Small direct edits could be a desirable alternative to repeating the command verbally. For a fallback, the user can type with a keyboard to replace the text. (Left: quick operation buttons for selecting pieces of the text)
(a) Finding Objects by Labels for Teleporting, Copying, Deleting
(b) Rule Toggling, Deleting
Figure 6. Find Panel: for performing search queries on sketches’ noun and adjective labels. (a) Selecting the word ”tree” searches for all objects with that label; tapping on an entry warps to the sketch; the eraser deletes the sketch instead; pen dragging copies entries (b) Forgetting active actions and rules: selecting the word ”actions” (in the transcript) and an object in the scene will display a panel of all current actions affecting the object. Deleting an action using the pen will stop the action immediately. Selecting the word ”rules” (in the transcript) will display active rules. Tapping with the pen will toggle the rule off/on.
4.4. Language Commands & Functionality
DrawTalking interprets the structure of language input into commands built from primitives. These primitives are intended to be a small sample of possible functionality that demonstrate our working concept. (They are partly inspired by existing software and design spaces (Shi et al., 2021; Dreams, 2020; Resnick et al., 2009).) (See the Info Sheet supplemental material.) We emphasize that our contribution is the interactive way of controlling elements. Our specific application and the fidelity and comprehensiveness of the supported behaviors, are implementation details independent of the DrawTalking concept.
Verbs are actions performable by sketches and the system, either built-in or user-defined in terms of other verbs by composition. Examples of implemented verb primitives include:
  • animations (e.g. move, follow, rotate, jump, flee),
  • state changes (e.g. create, transform),
  • events (e.g. collides with, press),
  • inequalities (e.g. equal, exceed).
Verb behavior changes based on other parts of the sentence:
Conjunctions:
run simultaneously (e.g. ”The dog jumps ”and” the cats jump”).
Sequences:
run in-order (e.g. ”The dog jumps ”and then” the cats jump).
Stop commands:
cancel an ongoing operation (e.g. ”The square stops moving”).
Prepositions:
(like on, to, under) cause verbs to exhibit different behavior, e.g. ”The dog jumps on/under the bed” impacts the dog’s final position relative to the target. All verbs can use any such prepositions as input. Prepositions also might describe the spatial relationship between objects in a command (Figure 7).
Timers:
specify the duration of a verb, e.g. ”the square moves up for 11.18 seconds and then jumps.”.
Loops:
repeat an action, either forever (e.g. ”endlessly the dog jumps”), or finitely (e.g. ”10 times the dog jumps excitedly”).
Special verbs include ”become”, which modifies the sketch’s labels, and ”transform into”, which also instantly replaces a sketch’s visual representation with another’s to support state changes, e.g. ”the sun transforms into a moon” or ”the frog transforms into a prince”. ”Follow” also offers a simple specialization on nouns — if a system object ”view” is commanded to follow a given object, the system camera will track that object. e.g. ”The view follows the hero.”
Nouns, pronouns, and deixis refer to object labels and are used to pick specific objects or specify types of objects. Using deixis while selecting an object will select the object immediately; pronouns can refer-back to objects, enabling more natural sentences as commands.
For choosing specific sketches in a command:
  • choose 1 or more specific object instances: ”The” $+$ noun
  • choose all of a label: ”all” $+$ nouns
  • select a specific number of objects: ¡number¿ $+$ noun(s)
  • select a random object: ”a” $+$ noun
There are a few special cases. The pronoun ”I” is reserved; it allows the user to take-part in the narrative of a command (e.g. ”I destroy the wall”). ”Thing” is also a reserved noun that can refer to any object regardless of label. Plural nouns with no modifiers (e.g. as in ”blades rotate”) refer to labels used to define the interactions between objects with that label. Specifically, this is used to construct rules, as described below and in Figure 8.
Adjectives and adverbs are usable as labels that define properties on objects. Verbs receive adjectives as arguments that are evaluated continuously to modulate values and effects (e.g. magnitude, speed, size, distance, height). For example, ”fast”, ”slow,” ”excited” impact jump height and/or movement speed. Adjectives also disambiguate like-noun labeled-objects (e.g. ”first house” vs. ”second house”). Adjectives can be removed programmatically via negation, e.g. The thing is not fast.
Adverbs heighten adjective effects multiplicatively – e.g. ”very” and ”slightly” – and are chainable – e.g. ”very, very.” Special adjectives offer system-control: e.g. ”static” fixes sketches to the screen like a GUI element, useful for buttons, d-pads, score displays, etc.; ”visible”/”invisible” toggle sketch visibility by command or automatically when saying ””This is a ¡…¿”.
Rules are conditionals that run any kind of command in the future when objects with certain labels satisfy the condition. This enables the user to specify automated commands for the future ”when,” ”as,” or ”after” an event has completed, without needing to know what they want or create objects in advance,
e.g., “When arrows collide with balloons, arrows destroy balloons.”. Rules also allow definition of new verbs in terms of existing primitives, e.g. ”flicker:” When lights flicker, forever lights disappear for 0.1 seconds and then lights appear for 0.1 seconds.
Figure 7. Hierarchical Disambiguation: Here, two sketches named ”blade” are attached to different parent objects, but the system picks the one with the correct relationship by looking at the names in the objects’ hierarchies. (The verb, ”attach to” was used to attach the sketches beforehand.)
4.5. Examples
Figure 8. Windmill Simulation. Flexible process for constructing an interactive windmill built from user-defined rules, sketches, and triggers. The user can do this in any order and can try results at each step. This works on any sketch labeled ”blade.”
[…]
[A mock-up showing the construction of a windmill whose blades rotate as wind sketches spawn-in and move past the blades, all scripted-in using speech commands for defining rules, along with hand-drawn sketches]A mock-up showing the construction of a windmill whose blades rotate as wind sketches spawn-in and move past the blades, all scripted-in using speech commands for defining rules, along with hand-drawn sketches; Panel 1 shows construction of the windmill with the base labeled ”windmill,” and the blades labeled ”blades.” The blades are attached to the windmill via the command (captioned below) ”I attach the blades to the windmill.” Panel 2 shows the user saving a hand-drawn wind spiral sketch so the system knows what to spawn when ”wind” is specified for a ”create” command. Panel 3 shows definition of rules for collision between wind and blades, captioned: ”When wind collides with blades, blades rotate.” and ”After wind collides with blades, blades stop rotating.” Panel 4 shows that extra ”switch” and left-side wall sketches have been drawn. Pressing the switch causes wind to spawn from the left wall and move to the right. Several clones of the previously saved wind sketch are shown moving to the right and colliding with the windmill blades, causing them to spin clockwise. Captioned are the commands to define all of this: ”When I press the switch I create wind at the wall.” ”After wind appears, wind moves right.”
Figure 9. Pong into Breakout. Left: per-player paddles and points; Middle, Right: scene reoriented, second player’s objects removed, destructible bricks logic defined and objects added in-bulk into the desired region. (Note: not all commands shown.)
[…]
[Two screenshots of (left) a game of Pong created in DrawTalking and (right) the same game’s sketches repurposed and extended into the similar game of Break-out]Two screenshots of (left) a game of Pong created in DrawTalking and (right) the same game’s sketches repurposed and extended into the similar game of Break-out; Left: Pong contains 4 wall sketches and score counters for each player atop the walls. There is a start game switch in the bottom right that starts the game when pressed. Two paddles within the walls are being manipulated via touch controls. A ball sketch is shown bouncing towards the right paddle, which is about to reflect it left. The current score is 0-0. Right: the scene has been rotated vertically, and the second score counter has been deleted. The user has created several blocks filling the top part of the inner-walled area of the game area they’ve created. Some of the blocks have been destroyed by the user’s bouncing ball. Currently, the ball has bounced off the user’s paddle and is heading towards another breakabout brick. The score counter reads 25 for the total number of points earned from breaking the blocks. All of these mechanics have been created from user-specified rules in speech commands.
Figure 10. Day/Night Cycle with Transforming Villagers. A sun/moon transforms into the other in-time with collision events. When the moon appears, the villagers transform into ghosts, which chase the boy. This shows event-based behavior changes.
Figure 11. ”Dog Plays Fetch” as a Prototype Game Mechanic (demonstrating adjectives for dynamic logic) The dog interaction is re-imagined as an interactive game mechanic.
We demonstrate DrawTalking with illustrated examples showing multiple procedural and programming-like capabilities.
4.5.1. Pond scene
A simple example demonstrating randomized behavior is shown in Figure 1. Here a user sketches a frog, lily pads, water, and a butterfly. To cause the frog to hop randomly between any existing or future lily pad, they say: ”Forever the frog hops to a lily,”. The butterfly is then commanded to follow the frog, and the user adjusts the speed of the butterfly: ”this butterfly is slow.”. The user can pause the simulation to edit any objects before resuming the actions.
4.5.2. Dog and boy’s infinite game of fetch
This highlights sequences, loops, and user participation. The user sketches a boy, dog, water, and ball and commands the boy and dog to interact with the ball (Figure 3). The user can play the role of puppeteer by interactively moving the objects (e.g. the ball) to influence the other sketches’ actions dynamically while the simulation is ongoing. For additional effect, the water could rise upon collision with the ball, with a command like: ”When water collides with balls water moves up for 0.2 seconds and then water moves down for 0.2 seconds.”
4.5.3. Windmill simulation
This example demonstrates rules, custom object saving and spawning, and buttons. (See Figure 8 for steps.) A key point is the flexibility to iterate on rules. In fact, here the rules can be defined in any order. If the user only defines the condition for on-collision, the blades will continue spinning by design. The user, however, can quickly rectify this by defining a rule to stop the blades after collision between wind and blades.
4.5.4. Creating the game ”Pong” and turning it into ”Breakout”
This example shows how one can quickly reuse functionality to turn one idea into another, in this case, a version of the game ”Pong” (noa, 2022) into a version of the game ”Breakout” (AtariAdmin, 2022) (Figure 9).
Create a ball and walls, then paddles, goals, and points for each player. For points, say ”I want the number 0” and touch+pen the number to spawn a number object. Label with ”This is the first/second score, goal.”. Make points screen-space UI with ”This thing is static” and set-up rules for collisions between the ball and goals: ”When balls collide with first/second goals second/first scores increase”. Last, add collision logic: ”when balls collide with walls walls reflect balls”, ”when balls collide with paddles paddles reflect balls.” Now we have a playable touch-based version of Pong.
We turn this into Breakout by rotating the canvas, deleting the second-player-related sketches, and adding a breakable blocks mechanic. To speed-up our process, we can sketch a temporary ”region” and say ”I pack the region with blocks” to fill it with sketches e.g. a custom ”block” sketch. To make the blocks destructible and increase the point count, create a rule ”When balls collide with blocks balls destroy blocks and then the score increases.” We are done. In a few steps we have revised the scene into a different game.
4.5.5. Day/Night Cycle: Transforming Objects and Periodic Collision Logic
The example in Figure 10 demonstrates the ability to command objects to transform into other saved objects automatically, which not only changes objects’ representation, but also enables automatic behavior and state changes. Further, the user can easily adjust periodic rates by using objects as a directly-manipulable, visual timing mechanism. By moving the walls in the figure closer or farther, we can opt to adjust the rate of transformation without the need for a precise timing value.
4.5.6. ”Dog plays fetch revisited” as a playable gameplay mechanic
The dog interaction is re-imagined as an interactive game mechanic. The example demonstrates adjectives being used for logical control (Figure 11). The user hits a button to trigger the dog’s fetch ability: collect a treat for points, but only if the treat is in the proximity of the boy’s collider. This example demonstrates a mix of logical behavior-programming using adjectives, as well as the beginnings of quick game-prototyping iteration.
4.6. Implementation
4.6.1. Hardware and Software
We implemented DrawTalking natively on the iPad Pro M2, with the main code written in C++ and C, bridged with Objective C++ and Swift to access platform-specific APIs. Speech recognition is on-device. A local NodeJS+Python server-side for NLP runs on a MacBook Pro. Text is continuously sent to the server where 1) a dependency-tree is created, 2) sent to the client and compiled into a key-value format for semantic roles (e.g. AGENT, OBJECT), 3) either interpreted in real-time by our custom engine to drive interactive simulations, or used to find objects by deixis. In final stages, the system looks-up and maps objects to roles in the command structure and executes it for animated effects. (See Figure 12 for a diagram of the the concrete client-server implementation and data-flow.)
4.6.2. Processing Steps
The language processing component comprises roughly 4 steps that translate a natural language data structure into a command for the application (Figure 13). For supplementary examples, see Appendix D.
Natural language raw text $\rightarrow$ S1
The system receives language input and outputs a directed graph data structure encoding the dependencies between and semantic roles of words.
Our implementation uses the Spacy (v 3.1.4)(Montani et al., 2023) library to output a dependency tree per-sentence, along with a library called ”coreferee” to fill-in coreference information.
S1 $\rightarrow$ S2
The system traverses S1 and generates a new generic graph structure S2. Each node entry in the graph structure represents a nested hierarchy of semantic units containing information such as labels (the word actually used) and part of speech (such as noun, verb, etc.).
Incomplete S2 $\rightarrow$ S2 with system-context feedback
Upon creating the structure for S2, there might be unspecified placeholders for objects, which we call ”incomplete”. The system will now look at application-context such as user-input and objects, and fill the structure with concrete entity IDs as-needed.
A query sub-system is needed to register and lookup objects based on their labels. For each noun-like entry in the traversal, the system queries for objects with the given noun and adjective labels, and returns the IDs of all objects in the world that match those labels.
Incomplete S2 $\rightarrow$ complete S2 with user feedback
After this process, the user should be able to do last-second editing and correction of the intermediate structure. The system can output user feedback (Figure 4).
S2 $\rightarrow$ S3
Lastly, the application traverses the structure and generates a final application-defined structure S3 that it can evaluate – in our case, a mix between scriptable virtual machine and animation engine akin to (Perlin and Goldberg, 1996; Resnick et al., 2009) or a runnable node-based program (cycling74, 2018). S3 can call these scripts and retain S2 for reference. When traversing nodes in S2 labeled ACTION (the verbs) we lookup the appropriate previously-defined script for that action and insert a reference to it in S3. The arguments for that action, i.e. semantic role keys mapped to object ids or types, are inserted into a lookup table specific to the ACTION script instance so when the action executes, it knows what objects to modify. Loops and timed waits are also inserted into S3. The application evaluates the final structure according to its own interpretation. This results in any potential application-specific side-effect – in our case, animation, simulation, rule creation, application state-changes and so on.
The application can also retain the generic S2 structure to generate commands later.
Figure 12. Concrete Client/Server Details. Text data are sent to a server for natural language processing over a data channel and forwarded to an NLP-focused sub-process on the same machine. The semantic parse data return to the client. A command channel handles user events independently (e.g. discarding commands interrupts in-flight stale NLP).
[…]
[A system diagram for the client-server architecture and data flow]A system diagram for the client-server architecture and data flow. The user is shown as a stick-figure with a quote bubble reading ”speech” and the text ”pen+touch” both connected to the main DrawTalking client node in the system diagram (indicating possible sources of user input.) The client node contains a real-life photo of the researcher’s iPad Pro used for the prototype, atop a blue square with the name ”DrawTalking” on it, as well as a small cartoon microphone icon. Below the square reads: ”Main C++, C; Platform: Obj C, Swift, Metal Graphics; On-Device Speech Recognition”. 3 arrows connect the client node to a server node. The arrows are referred to as ”Wireless Local Connection” (to the server. Arrow 1 (outgoing) is for text data. Arrow 2(incoming) is for the semantics parse. Arrow 3 is for control messages to the server; In the server node is a real-life photo of the Macbook Pro used for the server, above two nodes: left: the control server labeled NodeJS; right the Language Processing Sub-Process labeled Python 3. 2 arrows between server nodes (both surrounded by an enclosing rectangle): arrow 1: (outgoing) is for text. arrow 2: (incoming) is for the resulting semantics parse, which is sent back to the control server and back to the DrawTalking client along the other arrow incoming to the DrawTalking client.
Figure 13. Language processing into commands (S1 to S3): S1 is any implementation-specific NLP, e.g. a dependency tree; S2 stores generic semantic role labels created from S1; S3 is an app-specific engine that receives and interprets S2.
[…]
[Language processing data structures pipeline]Language processing data structures pipeline. 3 parts, S1, S2, S3: S1 ”Backend NLP Method of Choice” - accompanied by an abstract NLP dependency tree diagram. S2: Generic Format for Semantic Roles, accompanied by the following object: ” Action Agent thing_id = # Preposition Object thing_id = # ”; S3: Our Application-Specific Engine, accompanied by a text description and pseudocode for the engine code. The description reads: ”Parser script generated, instantiates script for action with S2 object lookup as input to act on objects specified in command Both scripts enqueued in Execution Engine. The pseudocode reads: ” // PSEUDOCODE // dynamically generates a custom script // for the command that runs every frame [¿(instruction index) [begin](wait Queue) // spawns a child script for action, // and adds to the local wait queue [call -¿ Action (with_args)] // parent blocks on “wait”, // checks completion status of // child script in queue [wait] [end] // script terminates ]
// dynamically instantiates the script // for the action, also runs every frame [ [begin](wait Queue) // script primitives specific // to the action go here … … [end] // sets completion status & terminates ] ”
5. Open-Ended User Study
We conducted an exploratory discussion-focused study to gauge understandability of interaction, discover use cases and directions, and learn users’ perceptions. We believe this form of verbal feedback is valuable because it can extend our own understanding of the tool with specific anecdotal and experiential feedback. The goal was to give the participant enough exposure to the interface to arrive at working artifacts from the process and elicit meaningful discussion.
To that end, the study was an open-ended exploration between the researcher and the participant. To mitigate the potential for biasing participants towards certain answers, the researcher did not reveal the design goals or indicate particular use cases for the tool. Participants were informed only that the tool was a prototype with animation features controlled by sketching and speaking. We let participants draw their own conclusions as to the usefulness, use cases, and potential for the interactions.
We chose this approach because we were interested in qualitative, early stage feedback on the concept and interactions behind DrawTalking, rather than feedback on our implementation, user performance, or based on comparisons. There is no known baseline for comparison with DrawTalking’s interactions, and at this stage, we felt it would not contribute towards achieving the goals of our research to compare the DrawTalking prototype with production tools in terms of usability (i.e. the ”usability trap” as in Olsen’s work (Olsen, 2007)). We wanted user feedback on potential use cases based on discussions rather than specializing on a specific use case.
5.1. Participants & Procedure
We invited 9 participants (students, professors, artists, game design) chosen by responses to an online recruitment form. We wanted candidates with relatively-high confidence in speaking and sketching, with interest in the topic, and with a fair mix of professional/academic backgrounds. For participant details, see Appendix B.
Each study lasted 1 hour and 15 minutes, and each participant was compensated by 30 USD. For each session, the researcher sat alongside the participant in front of DrawTalking running on an iPad, and taught the drawing features. Next, the participant was told to draw 5 objects of their choice, then taught both object labeling methods. Then, the session was improvised using the objects to explore all language features in rough increasing complexity. The researcher could help or suggest ideas, but participants mainly guided their own exploration using a provided language features list (the Info Sheet supplemental material), and each experience was different and ended with different interactive scenes. Participants were allowed to think-aloud and comment. After, the participant was asked to reflect on the experience. The researcher ended by showing additional pre-recorded videos and soliciting final feedback. We analyzed the results by taking note of common key words (e.g. fluidity) and of unique anecdotes, suggestions, and comments from the session and post-discussion. We recorded each session and took screen-shots for post-viewing (For samples, see subsection B.2). In the following section, we report the discussion-oriented results.
6. Results and Discussion
All participants understood and learned the mechanics, and produced thoughtful discussion. They tried their own ideas for how to make things or test their own understanding out of curiosity. All became comfortable with the core mechanics and some were excited to continue playing with the tool. They identified use cases include educational applications, rapid videogame and design prototyping, paper prototyping, user interface design, presentation, language learning, and visual-oriented programming.
We explore the qualitative experiences of participants using DrawTalking. The following come from the open-ended discussions and roughly covers topics related to the design goals.
6.1. As programming-like capability
DrawTalking’s style of control shows promise for empowering both programming and non-programming audiences (D4).
Several participants with programming knowledge and experience teaching programming compared DrawTalking functionality to constructs like variables, loops, and conditionals, indicating the patterns were familiar (D4). Some were reminded of Scratch (Maloney et al., 2010) if it did not have the explicit programming interface.
In contrast, P7 (an experienced digital artist without programming experience) appreciated DrawTalking as an accessible tool for non-programmers that could reduce frustration encountered with existing programming tools: ”Me instructing a game engine – someone like me who’s not a programmer and who is intimidated by doing C# or … using [a] visual scripting language like blueprints – this is a really clean interface that I think can achieve you can get 90- or 80% the way there. It just makes the user experience cleaner than having to use all these knobs and buttons or things like that or using scripting language or having to actually write code. You [the researcher] not force me to work with the scripting language. ”. P7 here favors the simplicity and accessibility of language control as a complement to more powerful programming interfaces.
The takeaway of interest is that not including an explicit coding interface was appealing for its simplicity and accessibility. Nevertheless, it would be interesting to explore balancing a greater level of complexity (in terms of interface) to serve more expert users, without compromising the simplicity for non-programmers. P7, for example, is aware of more powerful tools, but prefers something simpler. DrawTalking, in concept, seems to serve the role as abstraction layer between language- and code-based control. We could think about different modes of information-hiding depending on the task and audience.
6.2. Semantics-visuals correspondence is valued
Participants valued the visual correspondence between visual elements and language embedded in DrawTalking’s sketches and components as a strength in the design. (P5: ”Mixing visual elements with language helps me understand it. More of that [in future explorations] could be good… drawing connections between [things] visually as I’m speaking.”)
For the semantics diagram, participants understood it as a helpful means to introspect and correct the system’s understanding of user input. P2: ”[The diagram] is pretty important. Especially if you’re getting to used to the program, you might say something wrong and it interprets it in a different way and you can correct it.” P8: ”This is a really cool debug thing. I immediately understand what happens when I read it.” Additionally, the interface simplifies the input, making it easier to read: P7: ”[It’s] showing what it’s going to do. It’s taking the verbs and cutting all the other stuff out. The feedback suggests that the diagram is valuable for instilling confidence in the system, and that the interface achieved D3.
For the text view, we observed users (e.g. P1, P5) habitually clearing the transcript using ”discard” on their own. We asked why and some participants reported wanting to keep the interface clean. We did not foresee this, but this user behavior serves a dual-purpose: the user is aware of what they say, which means they are perhaps likelier to catch errors, and they also help by keeping the system input from growing too long. We can take this as a positive: the system helps the user, and the user helps the system. We can also take this as room for improvement: how to keep the text view easier to clean without user intervention?
6.3. Physicality, Prototyping, Playfulness
A reported strength of DrawTalking was the combination of physicality with logic and animation: P5: ”I can execute different rules of things that are happening. The scene is playing out here AND physical thing is happening.” P2, a games student, likened DrawTalking to a form of digital, semi-automatic paper-prototyping: ”When we make games there’s something called a paper prototype where we make a bunch of pieces of the players and the objects. We just move it around by hand to kind of simulate what it would be. [DrawTalking] is kind of like that but on steroids a bit. So it’s very nice to be able to kind of have those tools to help you with it without needing to manually make it.” Relatedly, P1 said, ”It’s like language and Legos put-together.”
This suggests that the design decision to give users interactive control via spatial direct manipulation in tandem with automation was successful; it enabled a form of playful exploration. This supports something akin to what professional designers and kids use to tinker with creative ideas. Note that P2 values ”help with making” rather than ”fully-automatic making,” meaning to them, interactivity was crucial for the playful prototyping feel.
6.4. Creative problem-solving
P1 was able to discover creative use of rule functionality by their own curiosity. In their case, they drew a boy, house, and tree. By creating two co-dependent rules, they wanted to see if they could recreate infinite loop functionality without using ”over and over” or ”forever.” Their solution was ”When boys collide with trees, boys move to houses” and ”When boys collide with houses, boys move to trees.”. This resulted in an infinite loop once the boy sketch collided with either the tree or the house. The example demonstrates the recombinability of the language-controllable primitives (Figure 16).
A highlight in P2’s session involved working-around a then-limitation of one of the verbs, ”climb.” The scene contained a tree and squirrel, and we wanted the squirrel only to climb the tree if it were not already on-top. However, our simple version of ”climb” was (at the time) programmed simply to move an object to the base of a target and then move it to the top, so the squirrel would inadvertently climb down then up. To work around this issue, a solution was devised in which an invisible collision box placed at the top of the tree programmed with ”When squirrels collide with the collider, squirrels stop climbing.”. Then, upon ”The squirrel climbs the tree”, the discovery was made that the collider could also be used as an in-situ stopping mechanism the user could simply drag onto the squirrel at any time to stop it (Figure 17).
We believe this means DrawTalking’s interactions successfully enable creativity, discoverability, and playful problem-
solving, independent of the limitations of the featureset.
6.5. Re-applicability to other toolsets
Users saw potential for DrawTalking as generic functionality for integration with other applications in a creative pipeline, e.g. as a playful ideation phase from which to import/export assets or runnable code; attaching speech control to production tools to support faster workflows. P1: ”You could sit here and have a conversation and build up just using language your interactions and then you send that out for [exporting] code.” P3: (comparing with an Adobe AfterEffects workflow) ”here, it just takes one second for me to ’say it’ – [this could be a] built-in function for any tool/interactive AI”. P3 felt the interactions were fluid. P8 emphasized language control: ”just incorporate the language and control interface here. We can easily create animations with the fancy stuff they have.”
In other words, users perceive DrawTalking as an interaction that is independent from a specific application. The suggestions for outputting code or plugging-into other applications with more advanced functionality hint at an interest in the way of controlling elements and a general excitement over how the ideas can be applied.
6.6. Naming by speech preferred to linking
We observed that labeling by speech (deixis) was unanimously preferred and used over linking and asked participants about their behavior. (See subsection 4.3 to recap labeling functionality.) It was considered more intuitive, direct, and similar to how people talk. On the other hand, pen+touch linking could be a useful fallback when freer grammar without directly-mentioned ”the” is preferred, e.g. when giving talks, or when referring to text far back in history without needing to speak again. Being able to link directly to a word could be more error-tolerant) P5 (an interactive computer graphics professor) gave such as example: ”I often talk about arrows in the context of vectors in my teaching”. Here, no deixis is used. Instead, the user would do a pen+touch operation between objects and the words ”arrows” and ”vectors” to do the labeling. This labeling can be performed in the background without requiring the speaker to use a specific grammar.
However, if the user does prefer the speech approach to the linking approach, this means it might be possible to reduce reliance on the transcript view, which could open-up opportunities to hide the transcript for contexts when it might be inconvenient. e.g. for eyes-free interfaces or XR spatial systems.
6.7. Greater control for physical interfaces
At least one participant, P9 was a roboticist who could give specialized feedback. When asked about possibly incorporating into robotics P9 cautioned that for real-world environments, they might want to have even more control and self-specification of commands due to safety concerns, in spite of the loss of automation:
P9: For potential tangible interfaces, physical things like robots might need more specification (for safety? accuracy?). [I] would be more tolerant of extra specification to make sure it’s correct. Definitely don’t want surprises. People won’t want it to be fully magical. [They want] more control.
This suggests that our emphasis on user control was necessary and perhaps crucial for re-application of DrawTalking concepts into future use cases working with physical objects for the sake of safety. There is a clear trade-off between ”feeling magical” and ”feeling in-control.”
7. Discussion & Future Work
In sum, users described DrawTalking as fluid and flexible; a natural language way of achieving programming-like functionality; a rapid prototyping environment; an independent general interaction technique; capable of integration with other applications; physical, tangible, spatial; accessible to kids; a new approach to working-out visual problems.
We believe DrawTalking works, then, because it successfully captures some of the playful and creative attributes of programming, spatial manipulation and sketching and language, owing to our initial goals and to our designs.
Based on the primary contribution of our work (the interaction mechanism enabling sketching+sketching control) concrete implementations could use the concept to drive deeper levels of functionality and programmability than what is capable with our application prototype. The overarching idea, above all, is about exploring how to provide controllable feedback and facilitate a kind of vernacular style of programming, to extend our creative capabilities.
There are many possible directions: Improving the system design and vocabulary; further-exploring visual-linguistic mappings; integrating with other applications; longer-term studies applying our approach to domain-specific creative or educational workflows; multi-user collaborations; application to spatial experiences for interaction with the real-world. We describe several future possibilities below:
Integration with other applications. A DrawTalking-like sketching application could represent a prototyping/open-ended phase of the creative process. Future work can explore how to take the interaction with these rough named sketches and convert them for use in later-stage production elements, where goals have become more specific. There is also an opportunity to explore how to convert back and forth between different representations of objects and behaviors to support moving seamlessly between ideation and production phases. For example, we suspect that our strategy of naming objects independent of their representation could be useful as a way of generating scenes from the sketch representation.
Semantics-visual mapping augmentation. There are several unexplored concepts relevant to semantics-visual mapping. For example, the use of sketches to represent abstract concepts could be evaluated as additional means for controlling content, in addition to the concrete sketches and language we’ve used. The semantics diagram, further, could be extended to support greater levels of editability. Another future area might involve context-sensitivity: What if the user’s culture and assumptions should produce different behaviors and visual-semantics mappings? How does DrawTalking, further, change when adapted to other languages?
Language input that is more natural, reducing user specification. This could potentially be addressed with a more robust and feature-complete back-end: e.g. improve the language processing implementation and/or integrate with faster generative language models for translation to simpler instructions.
Generative language models as of writing are too slow for interactive time, but in the future might be used to convert fully natural language into structured subsets easy for systems like ours to parse, or to generate commands based on application or scene context to reduce user-over-specification. This could allow for creation and management of larger persistent simulations.
Spontaneous authoring of new primitives without programming. Exploring interactions for richer behavior-authoring interactions compatible our approach would be an excellent direction. We can imagine a number of possibilities: exploring extensions to the existing semantics diagram interface for lightweight editing; using DrawTalking’s controls to drive other applications with their own unique primitives; combining with the capabilities of generative models; crowd-sourcing programmed-functionality through shared libraries of domain-specific functionality; introducing multiple layers of programming capability in the same interface for different audiences (e.g. artists, programmers), trading-off simplicity for granular control as is common in many applications.
Possibility for generalizability & scalability. We envision that combinations of DrawTalking-like functionality and different implementations and domain-specific applications could lead to a more-scalable ecosystem of interfaces. We also believe it may be useful to decouple the implementation of natural language processing from more deterministic processing, to facilitate reproducible research and support stabler interactive environments; in other words: stages of language compilation from probabilistic to deterministic instructions.
Lastly, although we’ve focused on rough sketching, conceptually there is no limitation preventing sketches from being replaced with other forms of controllable digital or tangible objects. A sketches can simply reduce to a collections of variable properties and representations.
We are excited to see ideas taken in new directions for playful and creative human-machine interfaces beyond our initial scope. We think that there’s an opportunity to seek-out more human-computer interactions that draw from our natural behavior for inspiration.
8. Conclusion
We have introduced DrawTalking, an approach to building interactive worlds by sketching and speaking, and have shown its potential. Our interface was inspired by our natural interplay between sketching, language, and our ability to communicate via make-believe. There are many exciting directions and we would be excited to see future research build on our approach or uncover other human-centered approaches to extend natural human abilities. We consider this project just one possible step and we hope that it will foster fruitful discussion and research.
Appendix A Formative Steps Additional Content
Figure 14 shows a sample of motivating content. Figure 15 shows P1’s result from the formative exercises.
Figure 14. Informal Formative Content Exploration: (top) subset of prepared and (bottom) live-performed sketch-based content for inspiration.
Figure 15. Formative Sketching Session, P1 Final Image. P1 narrated their own bird-watching story while sketching. Their iterative workflow and visual style of moving between ”islands of locations” helped inspire some features: e.g. discrete objects, object hierarchies, and spatially-flexible object relationships. Drawn roughly in-order: 1) Inline text titling the ”story”; 2) drawing of the city Toronto to set the location; 3) P1 visited the forest in Toronto; 4) P1 saw a bird, drew arrows indicating object relation (on a tree, in the forest); 5) P1 (the person sketch) found a pond and watched ducklings; 6) P1 took 200 photos.
Appendix B Exploratory Study with DrawTalking — Additional Content
B.1. Participants’ Background
The following summarizes the participants’ backgrounds:
P1expl:
An experienced professor (at a higher-level institution) in computer science, interactive graphics, and new visual media (e.g. AR, VR storytelling and games) — focuses on teaching students new to programming. Uses many visual aids in classes.
P2expl:
Digital fine arts university student (design and computer science), drawing is a hobby. Uses sketching for game design; planning gameplay features and figuring out how they should work. e.g. diagramming and framework design. Experienced with tablet drawing.
P3expl:
Experience as a digital designer, economics and digital art university student, not professionally-trained in art, uses a digital canvas (Procreate) to record visual ideas.
P4expl:
Tech / Design university student. Does design for web, desktop, apps in-general. Has done digital illustration and drawings often since childhood.
P5expl:
A very experienced professor in computer programming for interactive graphics. Does live-coding and streaming for internet-based education as well. Uses and authors open-source coding libraries for interactive graphics. Teaches programming for creating visuals rather than designing visuals by hand.
P6expl:
Student in interactive media. Draws and paints using the physical medium, with lifelong interest and experience. Experienced with digital graphical design work. Low experience with programming.
P7expl:
Industry creative expert. 15+ years experience as a set designer and in the theater industry. Digital illustrator, spatial experience designer, AR/VR immersive projects, projects in entertainment and fine-arts. Creative directing, collaborations with companies and academia. Non-programmer, some experience with visual blocks-based interfaces.
P8expl:
Robotics and machine-learning researcher, works with developing ML models, often needs to visualize ML inputs/outputs and illustrates rough sketches to think-through policies (for robots). No prior experience using a tablet.
P9expl:
Robotics and machine-learning PhD student, used tablet-based sketching interfaces.
B.2. Select Screenshots
(a) Define: upon collision with trees, boys move to houses.
(b) Define: upon collision with houses, boys move to trees.
Figure 16. P1: Creatively defining a custom loop using rules. P1 discovered a way to make a kind of infinite loop using 2 rules to cause an infinite sequence of collision/response movements back and forth between the house and the tree.
Figure 17. P2: Squirrel Climbing Tree Stopped by Collider. A rule is assigned to the collider to make it a prop the user can move around themselves to stop the squirrel from climbing upon collision with the collider.(The collider is shown in the semantics diagram as a rectangle, but invisible in the canvas on-purpose — where the label ”collider” is.)
Figure 18. P3: ”When seeds appear birds follow seeds.”
Figure 19. P5: ”The cat jumps on the mountain two times.”
Figure 20. P6: ”Pluto revolves around the sun.”
Figure 21. P7: ”The snake follows the bird.”
Figure 22. P8: Space scene with robots, ducks, asteroids, and stars near Mars.
Figure 23. P9: ”The girl jumps on the comfy chair.”
A selection of examples from the user sessions are shown: P1 Figure 16, P2 Figure 17, P3 Figure 18, P4 Figure 2, P5 Figure 19, P6 Figure 20, P7 Figure 21, P8 Figure 22, P9 Figure 23.
Appendix C Exploratory Features and Examples for the Record
During the development of this project, we explored several ideas mixing interactive visuals and text. Some of these did not contribute to the core ideas and this leg of the research was complete without them, but the examples might be useful to include for readers interested in combining more direct-manipulation techniques.
For example: reusable text commands in Figure 24 and context-sensitive representations of numerical outputs based on labels in Figure 25.
Figure 24. Text objects as macros. We can reduce repetitive speech by storing text as inline objects that work exactly as text in the semantics diagram (Figure 4). They’re created by linking the text transcript’s current content and the canvas. Here, inline text objects (in purple) have been created from 1) the word ”car” and 2) the sentence, ”the car moves right for five seconds.” The user can link a new object with 1) to label it ”car” and link 2) and the language action button to treat the text as a reusable command. (Running ”the car moves right…” again.) There is potential to explore this further. However, we decided to focus on the more general capability of triggering commands with rules, e.g. upon pressing a button.
Figure 25. Context-sensitive representations based on labels: We briefly explored using the labels to control the visual representation of number sketches. In this case, arrows connected from a freehand sketch output the sketch’s angle to the numbers, and the labels determine whether to display the angle in degrees or radians. ”Dynamic representations” was left to future work.
We also prototyped an additional working example specializing in ”digital interactive assignments” in which the user builds a molecule-matching game. See Figure 26. An empty ”box” sketch, a ”checkbox” sketch, and a hierarchical ”water” sketch have been saved prior that looks like the structure shown, with ”water” labeled. (These were deleted prior.) Next, the rule was created: ”When atoms form water, the box transforms into a checkbox.”. This creates a script that checks whether the user has created any sketch structure matching the names and hierarchy of the target. We believe the matching functionality of ”form” would be useful to have available in an interactive lesson notebook or textbook.
Figure 26. Molecule Matching Lesson Prototype - This example shows a checkbox (initially empty: a) that has automatically been filled (b) after the user has recreated the structure of a sketched water ($H_{2}O$) molecule formed from sketches labeled atoms, hydrogen, and oxygen.
Appendix D Abstract Semantic Structure
The following defines the basic structure of S2 (subsubsection 4.6.2), which is a simplified semantic structure graph representing an interpretable command. The concrete implementation of DrawTalking traverses this structure to generate final execution commands S3 (subsubsection 4.6.2).
Listing 1: S2 Data Structure
[…]
type :=
CMD_LIST |
ACTION |
AGENT | DIRECT_OBJECT | OBJECT | INDIRECT_OBJECT |
PREPOSITION |
TRIGGER_RESPONSE | TRIGGER | RESPONSE |
SEQUENCE_SIMULTANEOUS | SEQUENCE_THEN |
PROPERTY | PLURAL | COUNT | SPECIFIC_OR_UNSPECIFIC | TIME |
COREFERENCE
value :=
Number | // e.g. any of Float64, Float32, Uint64, etc.
Thing_ID |
Thing_Type |
String |
Boolean |
// e.g. pointer or int ID to dynamic-allocated object
Reference |
List[Value_Type]
// S2_Elements should be allocated with stable pointers, or use stable IDs
parent : Reference(S2_Element)
// Similar to JSON, but elements are always lists
// (can have multiple children for the same key
// (although the layout is not a hard requirement)
key_to_value := Map[String : List[S2_Element]]
// property can refer to another property e.g. for coreference
refers_to : Reference(S2_Element)
// usually refers to some user feedback UI element
user_feedback_ref : Reference(Anything)
// optional: stable pointer or ID to the raw token in
// the language input used to create this element
token : Reference(Token)
Listing 2: Possible structures for S2
// Note that each right-hand-side can be a list.
CMD_LIST -> any of the rest
ACTION -> any combination of
SOURCE,
DIRECT_OBJECT,
INDIRECT_OBJECT,
OBJECT,
PREPOSITION,
// contains a "trait" with a string value for the property name
// e.g. how adverbs or adjectives are used.
PROPERTY,
SEQUENCE_SIMULTANEOUS,
SEQUENCE_THEN,
// usually how long the action should last
TIME
COREFERENCE,
TRIGGER_RESPONSE -> TRIGGER + RESPONSE
TRIGGER, RESPONSE -> ACTION
// equivalent to ACTION.
// (but handled differently to generate rules.
// TRIGGER should be used to generate rules.
// RESPONSE should be used to generate commands
// invoked in the future with arguments generated from rule evaluation).
PREPOSITION -> OBJECT
SOURCE, DIRECT_OBJECT, INDIRECT_OBJECT, OBJECT -> any combination of
PLURAL, // whether plural or not
COUNT, // number of elements
SPECIFIC_OR_UNSPECIFIC // referring to a specific object or not
PROPERTY
PLURAL, COUNT, SPECIFIC_OR_UNSPECIFIC, TIME are terminal
SEQUENCE_THEN -> ACTION
SEQUENCE_SIMULTANEOUS -> ACTION
COREFERENCE // contains a pointer to another node in the value, usually noun-like.
Listing 3: Output from ”Forever the person throws the ball into the pond and then the dog gives the ball to her.” prior to object selection
label=[], tag=[], type=[], kind=[], key=[], idx=[0] id=[1931]
[CMD_LIST] = [
label=[], tag=[], type=[CMD], kind=[], key=[CMD_LIST], idx=[0] id=[1932]
[ACTION] = [
label=[throw], tag=[VERB], type=[ACTION], kind=[ACTION], key=[ACTION], idx=[0] id=[1933]
[PREPOSITION] = [
label=[into], tag=[], type=[into], kind=[], key=[PREPOSITION], idx=[0] id=[1934]
[OBJECT] = [
label=[pond], tag=[NOUN], type=[], kind=[THING_INSTANCE], key=[OBJECT], idx=[0] id=[1935]
value={
THING_INSTANCE=[609]
[SPECIFIC_OR_UNSPECIFIC] = [
label=[the], tag=[DET], type=[VALUE], kind=[SPECIFIC], key=[SPECIFIC_OR_UNSPECIFIC], idx=[0] id=[1936]
value={
FLAG=[true]
[COUNT] = [
label=[], tag=[], type=[VALUE], kind=[], key=[COUNT], idx=[0] id=[1937]
value={
NUMERIC=[1.000000]
[PLURAL] = [
label=[], tag=[], type=[VALUE], kind=[], key=[PLURAL], idx=[0] id=[1938]
value={
FLAG=[false]
[SEQUENCE_THEN] = [
label=[give], tag=[VERB], type=[ACTION], kind=[ACTION], key=[SEQUENCE_THEN], idx=[0], @=[MUST_FILL_IN_AGENT] id=[1939]
[PREPOSITION] = [
label=[to], tag=[], type=[to], kind=[], key=[PREPOSITION], idx=[0] id=[1940]
[OBJECT] = [
label=[person], tag=[NOUN], type=[], kind=[THING_INSTANCE], key=[AGENT], idx=[0] id=[1968]
<coreference substitution> id=[1960]
value={
THING_INSTANCE=[606]
[SPECIFIC_OR_UNSPECIFIC] = [
label=[the], tag=[DET], type=[VALUE], kind=[SPECIFIC], key=[SPECIFIC_OR_UNSPECIFIC], idx=[0] id=[1969]
value={
FLAG=[true]
[COUNT] = [
label=[], tag=[], type=[VALUE], kind=[], key=[COUNT], idx=[0] id=[1970]
value={
NUMERIC=[1.000000]
[PLURAL] = [
label=[], tag=[], type=[VALUE], kind=[], key=[PLURAL], idx=[0] id=[1971]
value={
FLAG=[false]
[DIRECT_OBJECT] = [
label=[ball], tag=[NOUN], type=[], kind=[THING_INSTANCE], key=[DIRECT_OBJECT], idx=[0] id=[1964]
<coreference substitution> id=[1955]
value={
THING_INSTANCE=[607]
[SPECIFIC_OR_UNSPECIFIC] = [
label=[the], tag=[DET], type=[VALUE], kind=[SPECIFIC], key=[SPECIFIC_OR_UNSPECIFIC], idx=[0] id=[1965]
value={
FLAG=[true]
[COUNT] = [
label=[], tag=[], type=[VALUE], kind=[], key=[COUNT], idx=[0] id=[1966]
value={
NUMERIC=[1.000000]
[PLURAL] = [
label=[], tag=[], type=[VALUE], kind=[], key=[PLURAL], idx=[0] id=[1967]
value={
FLAG=[false]
[AGENT] = [
label=[dog], tag=[NOUN], type=[], kind=[THING_INSTANCE], key=[AGENT], idx=[0] id=[1951]
value={
THING_INSTANCE=[608]
[SPECIFIC_OR_UNSPECIFIC] = [
label=[the], tag=[DET], type=[VALUE], kind=[SPECIFIC], key=[SPECIFIC_OR_UNSPECIFIC], idx=[0] id=[1952]
value={
FLAG=[true]
[COUNT] = [
label=[], tag=[], type=[VALUE], kind=[], key=[COUNT], idx=[0] id=[1953]
value={
NUMERIC=[1.000000]
[PLURAL] = [
label=[], tag=[], type=[VALUE], kind=[], key=[PLURAL], idx=[0] id=[1954]
value={
FLAG=[false]
[DIRECT_OBJECT] = [
label=[ball], tag=[NOUN], type=[], kind=[THING_INSTANCE], key=[DIRECT_OBJECT], idx=[0] id=[1955]
value={
THING_INSTANCE=[607]
[SPECIFIC_OR_UNSPECIFIC] = [
label=[the], tag=[DET], type=[VALUE], kind=[SPECIFIC], key=[SPECIFIC_OR_UNSPECIFIC], idx=[0] id=[1956]
value={
FLAG=[true]
[COUNT] = [
label=[], tag=[], type=[VALUE], kind=[], key=[COUNT], idx=[0] id=[1957]
value={
NUMERIC=[1.000000]
[PLURAL] = [
label=[], tag=[], type=[VALUE], kind=[], key=[PLURAL], idx=[0] id=[1958]
value={
FLAG=[false]
[PROPERTY] = [
label=[modifier], tag=[ADV], type=[PROPERTY], kind=[], key=[PROPERTY], idx=[0] id=[1959]
value={
TEXT=[forever]
[AGENT] = [
label=[person], tag=[NOUN], type=[], kind=[THING_INSTANCE], key=[AGENT], idx=[0] id=[1960]
value={
THING_INSTANCE=[606]
[SPECIFIC_OR_UNSPECIFIC] = [
label=[the], tag=[DET], type=[VALUE], kind=[SPECIFIC], key=[SPECIFIC_OR_UNSPECIFIC], idx=[0] id=[1961]
value={
FLAG=[true]
[COUNT] = [
label=[], tag=[], type=[VALUE], kind=[], key=[COUNT], idx=[0] id=[1962]
value={
NUMERIC=[1.000000]
[PLURAL] = [
label=[], tag=[], type=[VALUE], kind=[], key=[PLURAL], idx=[0] id=[1963]
value={
FLAG=[false]
Listing 4: Output from ”Every few seconds the frog hops to a lily.” prior to object selection
label=[], tag=[], type=[], kind=[], key=[], idx=[0] id=[1018]
[CMD_LIST] = [
label=[], tag=[], type=[CMD], kind=[], key=[CMD_LIST], idx=[0] id=[1019]
[ACTION] = [
label=[hop], tag=[VERB], type=[ACTION], kind=[ACTION], key=[ACTION], idx=[0] id=[1005]
[PREPOSITION] = [
label=[to], tag=[], type=[to], kind=[], key=[PREPOSITION], idx=[0] id=[1013]
[OBJECT] = [
label=[lily], tag=[NOUN], type=[], kind=[THING_INSTANCE], key=[OBJECT], idx=[0] id=[1014]
value={
THING_INSTANCE=[0]
[SPECIFIC_OR_UNSPECIFIC] = [
label=[a], tag=[DET], type=[VALUE], kind=[UNSPECIFIC], key=[SPECIFIC_OR_UNSPECIFIC], idx=[0] id=[1016]
value={
FLAG=[false]
[COUNT] = [
label=[], tag=[], type=[], kind=[], key=[COUNT], idx=[0] id=[1017]
value={
NUMERIC=[1.000000]
[PLURAL] = [
label=[], tag=[], type=[VALUE], kind=[], key=[PLURAL], idx=[0] id=[1015]
value={
FLAG=[false]
[TIME] = [
label=[second], tag=[TIME], type=[INTERVAL], kind=[], key=[TIME], idx=[0] id=[1006]
[PROPERTY] = [
label=[trait], tag=[ADJ], type=[PROPERTY], kind=[], key=[PROPERTY], idx=[0] id=[1008]
value={
TEXT=[few]
[AGENT] = [
label=[frog], tag=[NOUN], type=[], kind=[THING_INSTANCE], key=[AGENT], idx=[0] id=[1009]
value={
THING_INSTANCE=[0]
[SPECIFIC_OR_UNSPECIFIC] = [
label=[the], tag=[DET], type=[VALUE], kind=[SPECIFIC], key=[SPECIFIC_OR_UNSPECIFIC], idx=[0] id=[1011]
value={
FLAG=[true]
[COUNT] = [
label=[], tag=[], type=[VALUE], kind=[], key=[COUNT], idx=[0] id=[1012]
value={
NUMERIC=[1.000000]
[PLURAL] = [
label=[], tag=[], type=[VALUE], kind=[], key=[PLURAL], idx=[0] id=[1010]
value={
FLAG=[false]